rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#8622 — FS#12490 — mtl-1-6k

Attached to Project— Network
Maintenance
Beauharnois
CLOSED
100%
Link flap between BHS and MTL, we are going to replace the optics on BHS side.
Date:  Thursday, 19 February 2015, 16:53PM
Reason for closing:  Done
Comment by OVH - Tuesday, 27 January 2015, 14:41PM

We are going to proceed with replacement of the concerned optics.
We are monitoring the CRC and flaps on the interface in order to close this task.


Comment by OVH - Tuesday, 27 January 2015, 15:09PM

The link flap is again present , we will work on the
mtl site to clean the lens ans optics .


Comment by OVH - Wednesday, 28 January 2015, 10:28AM

We continue to monitor the ports in question.


Comment by OVH - Monday, 02 February 2015, 20:53PM

Again, flaps appeared. We will replace the entire optical path on MTL, the same action had no result on BHS.


Comment by OVH - Wednesday, 04 February 2015, 16:15PM

An optic cable has been replaced at MTL. If this doesn't fix the issue, we will change all the jumpers in this area.


Comment by OVH - Thursday, 05 February 2015, 10:28AM

The optic cable has been replaced in MTL, but the connection is still down. We'll change the jumpers asap.


Comment by OVH - Monday, 09 February 2015, 13:34PM

We've replaced the jumper and the connection is back up. We will monitor it.


Comment by OVH - Thursday, 19 February 2015, 16:53PM

The flap is back again.

We've found a faulty connection on a patch. The flap is patched. We will continue to monitor it.